# A Large-Scale Privacy Assessment of Android Third-Party SDKs

### v1

@@ Abstract.
Third-party Software Development Kits (SDKs) are widely adopted in Android app development, to effortlessly accelerate development pipelines and enhance app functionality. However, this convenience raises substantial concerns about unauthorized access to users’ privacy-sensitive information, which could be further abused for illegitimate purposes like user tracking or monetization. Our study offers a targeted analysis of user privacy protection among Android third-party SDKs, filling a critical gap in the Android software supply chain. It focuses on two aspects of their privacy practices, including data exfiltration and behavior-policy compliance (or privacy compliance), utilizing techniques of taint analysis and large language models. It covers 158 widely-used SDKs from two key SDK release platforms, the official one and a large alternative one. From them, we identified 338 instances of privacy data exfiltration. On privacy compliance, our study reveals that more than 30% of the examined SDKs fail to provide a privacy policy to disclose their data handling practices. Among those that provide privacy policies, 37% of them over-collect user data, and 88% falsely claim access to sensitive data. We revisit the latest versions of the SDKs after 12 months. Our analysis demonstrates a persistent lack of improvement in these concerning trends. Based on our findings, we propose three actionable recommendations to mitigate the privacy leakage risks and enhance privacy protection for Android users. Our research not only serves as an urgent call for industry attention but also provides crucial insights for future regulatory interventions.
1. Introduction
@@
The advent and ubiquity of third-party Software Development Kits (SDKs) in Android app development herald a profound shift in the mobile software development paradigm, transforming efficiency and accelerating the pace of innovation. These SDKs provide pre-built reusable code that offers a broad spectrum of functionalities, from fundamental features like user authentication (authe_lib, ) and data encryption (encryption_lib, ) to specialized components like analytics (analytics_lib, ), advertisement delivery (advertisement_lib, ), and user interface enhancements (UI_lib, ). The open-source nature and global reach of the Android ecosystem further fuel the proliferation of these SDKs.
@@
Integrating an SDK into an app is often as simple as declaring dependencies in the build configuration file, thus drastically reducing development time, and allowing developers to focus on app-specific features. Nowadays, the majority of Android apps incorporate multiple third-party SDKs, with an average app integrating approximately 8.6 SDKs (salza2018, ; zhang2018detecting, ). However, this widespread dependency also introduces significant security and privacy concerns. The complexity of the software supply chain and the opaque data handling practices of some SDKs can lead to privacy risks and security vulnerabilities, either due to malicious intent or unintentional oversights. These can be exploited by malevolent actors or lead to inadvertent privacy violations, as evidenced by recent high-profile incidents related to third-party SDKs (priv_leak_blackkite, ; priv_leak_mahajan, ).
@@
The privacy issues caused by SDKs could be further magnified in the pre-installed apps on Android devices. These apps are bundled by device manufacturers or network carriers, and therefore, often have an extremely great user population. They come with high privileges and are typically non-removable, such that an SDK embedded by them can stealthily perform higher-privileged operations that normal apps are unable to. An illustrative example is an app named “Mobile Services Manager” pre-installed on certain Android devices, which was found to download apps without user consent due to a malicious third-party SDK embedded (htfma2023msm, ; mollah2022mobile, ).
@@
Assessing privacy practices of SDKs includes two key aspects, i.e., data exfiltration and behavior-policy compliance (or privacy compliance). The data exfiltration involves the behaviors of locating sensitive data in publicly accessible places, such as the publicly accessible directories in file systems and the Internet. The opacity of SDKs’ operations and their “black-box” nature hinders meaningful assessments, leaving both end-users and app developers in the dark and elevating risks. Dishonest or malicious SDKs may disguise as their embedding apps to extensively collect privacy-sensitive data, while the end-users are unable to distinguish which party requests their privacy data. Incidents like the X-Mode controversy (keegan2022sold, ), where user locations were tracked by over 100 apps through embedded SDKs, underscore this point and serve as a stark reminder of the challenges associated with preserving user’s privacy in third-party SDKs.
@@
On the regulatory front, data protection regulations have been put in place globally, such as the General Data Protection Regulation (GDPR) (gdpr2016, ) and the California Consumer Privacy Act (CCPA) (ccpa2023california, ), which impose a stringent requirement on data processors. In response to this, software developers are required to provide a privacy policy to explicitly disclose what personal information is collected and how these data are handled by their software. Dissecting whether the data handling behaviors of SDKs comply with the claims in their privacy policies, which we define as behavior-policy compliance (or privacy compliance), becomes an urgent task.
@@
In this work, we conduct a large-scale study on the privacy practices of existing third-party SDKs. Our study covers 158 prevalently-installed SDKs from two mainstream release platforms, including Google Play SDK Index (google_play_sdk_index, ) and CAICT SDK Index (caict, ). The former is the official source of Android SDKs as a part of Google Play infrastructure, which lists the most widely used commercial SDKs in the Android app ecosystem (rodriguez2024sharing, ). The latter is claimed to be the largest release platform of mobile SDKs in China, serving as a complementary source of SDKs due to the absence of Google’s service in the Chinese market. Based on the statistical data provided by the two platforms, over 100 collected SDKs have been embedded by at least 100 apps with over 10 million installations, demonstrating the broad impact of this study.
Our study comprises two main strategies, i.e., conducting a data exfiltration assessment through static program analysis, and assessing privacy compliance through interpreting privacy policies. Specifically, we conduct a static taint analysis for tracking the flow of privacy-sensitive data within the SDK. We explore the traces that read and share privacy-sensitive data in public spaces, and accordingly investigate whether these traces can lead to potential privacy-sensitive data exfiltration. For the privacy compliance assessment, we employ a large language model (LLM) to carry out a text analysis of the SDKs’ privacy policies to identify what data is requested, cross-check the data claimed in privacy policies with the data access behaviors caught during the taint analysis, and determine if an excessive collection or over-claiming issue exists.
@@
  • Detecting Potential Privacy Data Exfiltration on a Large Scale. We collect 158 SDKs from two mainstream repositories and leverage taint analysis to track sensitive data flow, thereby exploring their potential data exfiltration. We managed to find 338 data exfiltration evidenced by sharing obtained privacy data into public spaces. Our work highlights the substantial privacy risks stemming from the widespread use of third-party SDKs in the Android ecosystem, unveiling complex and often opaque data-handling practices.
@@
2. Background and Problem Definitions
2.1. Android SDKs and Data Collection Practices
Android third-party SDKs represent a crucial component within the broader ecosystem of Android app development. These software SDKs offer pre-written code to app developers, enabling the efficient integration of a myriad of features and functions within their applications, ranging from user interface elements to complex data analytics capabilities. By leveraging these SDKs, developers can expedite the application development process, avoiding the necessity to “reinvent the wheel” for common functionalities. Consequently, these SDKs have become an integral part of the Android development framework, enhancing productivity and fostering innovation through code reuse.
Data collection practices associated with Android third-party SDKs, however, have been a subject of scrutiny within academic and industry communities. It has been observed that some third-party SDKs, while providing utility functions, also engage in extensive data collection activities without explicit acknowledgment or consent received from the end-users. These practices tend to infringe upon user privacy at times, collecting sensitive information such as device identifiers, location data, and personal preferences. Furthermore, there are instances where data is transmitted to remote servers, raising concerns about data security and potential misuse. In light of this, there is an increasing emphasis on implementing robust privacy and security measures, such as privacy-preserving APIs and data anonymization techniques, within these SDKs to ensure ethical and lawful data collection practices.
2.2. Objectives and Problem Definitions
This work aims to (1) detect potential exfiltration of users’ privacy data existing in the third-party SDKs, (2) explore the privacy compliance of developers in the SDK release, and thereby, (3) unveil the landscape of privacy protection in the Android ecosystem at the SDK level. To this end, we propose a compliance model to systematically assess the privacy practice of Android SDK regarding personal information collection. We let $\mathbb{D}$ represent a set of diverse data considered as user privacy, and let $\mathbb{S}$ be the set of publicly released third-party SDKs. Then we define what privacy data an SDK claims to collect and actually collects as follows.
@@ 2.2. Objectives and Problem Definitions
We let $\mathcal{C}$ denote the compliance disclosure of an SDK $\mathbf{s}\in\mathbb{S}$. Specifically, it defines the set of privacy data claimed to access in the privacy policy of the SDK $\mathbf{s}$. It shall also be assumed not to leak any privacy data to the public.
@@
The compliance practice of an SDK concerns two types of operations in its implementation, i.e., collecting privacy data, and sharing the collected privacy data. We let $\mathcal{R}_{s}$ denote the scope of privacy data would be read/collected by an SDK $\mathbf{s}$ based on its implementation, in which $\mathcal{R}_{s}\subseteq\mathbb{D}$. We also let $\mathcal{U}_{s}$ denote the collected privacy data involved in any uploading/sharing operations, given that $\mathcal{U}_{s}\subseteq\mathcal{R}_{s}$. Thus, we have the compliance practice of an SDK $\mathbf{s}$ as a tuple of its collecting and sharing behaviors, written as $\mathcal{P}_{s}=(\mathcal{R}_{s},\mathcal{U}_{s})$.
@@
Sharing users’ privacy in public spaces poses incompliance risks especially when the collection or the sharing operations are not performed with users’ consent. Moreover, some regulations like the GDPR have explicitly prohibited any collection of UUIs, amplifying the risks of privacy sharing on the SDK side. Here we define that any pre-recognized privacy data $\mathbf{d}$ being observed in sharing operations of an SDK $\mathbf{s}$ would constitute privacy sharing risk, detailed as follows:
(1) $\exists s\in\mathbb{S},\exists d\in\mathbb{D},d\in\mathcal{U}_{s}\Rightarrow\,\,\nvdash_{\mathcal{P}}\mathcal{C}_{s}.$
@@
(2) $\exists s\in\mathbb{S},\exists d\in\mathbb{D},(d\in\mathcal{R}_{s})\land(d\notin\mathcal{C}_{s})\Rightarrow\mathcal{R}_{s}\nvdash_{\mathcal{P}}\mathcal{C}_{s}.$
We remark that Google has not set a formal definition of “data collection” at the SDK level as what it does for third-party apps (google2024understanding, ). Thus, we refer to the guideline of GDPR Article 13 (gdpr2016, ; wolford2024guide, ) and take a conservative strategy by considering data accessed by SDKs as collected when evaluating the excessive collection, given that the data accessed by an SDK can flow into the host apps, which can then consume or send the data out.
@@
We assert an SDK that has an over-claiming issue if it is found to claim its access to more types of privacy data than what it actually reads. Over-claiming issues, although may not substantially lead to privacy data exfiltration, seriously violate the data minimization principle stipulated in Article 5(1)(c) of GDPR. It is defined as follows.
(3) $\exists s\in\mathbb{S},\exists d\in\mathbb{D},(d\notin\mathcal{R}_{s})\land(d\in\mathcal{C}_{s})\Rightarrow\mathcal{R}_{s}\nvDash_{\mathcal{P}}\mathcal{C}_{s}.$
2.3. Threat Model
We consider the exfiltration of users’ privacy data caused by malicious SDKs should follow a unified routine. To achieve such exfiltration, the malicious SDK (the attacker) gains capabilities as follows:
@@ 2.3. Threat Model
3. Scope of Privacy Data
Table 1. List of 41 pre-identified privacy data types and the distribution of the 741 source methods
@@ 3. Scope of Privacy Data
Figure 1. The workflow of potential data exfiltration exploration and compliance assessment at the SDK level
Sources
We resort to the Android developer documentation, including the permissions and OS release reports, to define the list of privacy data. Specifically, we browse the latest list of permissions defined in Android OS (google2024manifest, ) to retrieve the data that request explicit permissions to access, and from which we identify privacy data, such as users’ location, contacts, and calendars. These privacy data are complemented with privacy-sensitive items recognized from the privacy or security change section of each Android OS release report, for example, the clipboard and a group of user-unresettable identifiers (UUIs) are recognized in the release document of Android 10 (google2023privacy10, ) and added to the scope of this study. In addition, we also take existing literature in personal information collection (chen2014information, ; leung2016should, ; papadopoulos2017long, ; ren2018longitudinal, ; ren2016recon, ; wang2021understanding, ) and app analysis (enck2014taintdroid, ; gibler2012androidleaks, ; qiu2015apptrace, ; sun2016taintart, ) into consideration to complement the list of privacy data. We thus include the usage of diverse sensors, including motion, light, and pressure perception, into the list of privacy data because they have been widely studied in the research community. As a result, we managed to recognize 41 types of privacy data. These 41 types of privacy data constitute $\mathbb{D}$ defined in Section 2.2, formulating the scope of our privacy compliance assessment. The list of recognized privacy data is shown in Table 1.
@@
  • We then refine our categorization by grouping those data that can be programmably accessed through APIs in the same or similar classes based on the Android developer’s documentation. This is to facilitate our next step’s taint analysis as we need to collect as many involved APIs as possible to be the source methods, which we will detail in Section 4.4.
@@
Based on the criterion listed above, we further categorize them into five classes, namely chip, cellular & peripheral (C1), wireless communication (C2), camera, location, and other sensors, (C3), media and software specific (C4), and miscellaneous personal data (C5). More details of the five categories can be found in the Appendix A.1.
@@
Considering the request of personal data in privacy policies may not always written in a precise manner as defined in the developer’s documentation, and sometimes certain privacy data items have multiple names or colloquial phrases, we refer to relevant research (andow2020actions, ; bui2021consistency, ) and propose an ontology to capture the semantic relationship among the 41 types of privacy data, and consequently to guide our assessment.
@@
Besides that, we manually collect alternative names or colloquial phrases for each type of identified privacy data and treat them as synonyms ($\equiv$). For example, we find that the advertising ID and GAID are interchangeably used in privacy policy documents to refer to the Google Ad ID. Thus, all the synonymous terms of a privacy data type will be considered in our assessment.
4. Methodology
4.1. Approach Overview
We propose a four-phase approach and briefly illustrate it in Figure 1. Our study begins with collecting as many Android SDKs as possible from the two mainstream release platforms, shown as Phase 1. We first retrieve all SDKs available from the Google Play SDK Index (google_play_sdk_index, ), which is the official source of SDKs according to Android developer documentation. In addition, we also resort to the CAICT SDK index (caict, ) as the supplement source due to the absence of Google Play service in the Chinese market. As a result, we have collected 158 SDKs. For each SDK, we collect its binary file, the privacy policy or relevant disclosure, and meta information from the product webpage of the release platform. We detail our approach for this phase in Section 4.2.
After collecting the SDKs, we analyze the privacy policies of the collected SDKs, from which we identify data requested by the SDKs (shown as Phase 2). Due to the varying writing quality and styles of SDK documents, conventional privacy policy analysis approaches based on semantic ontologies and specific patterns or keywords cannot exhibit the desired performance. For that reason, we resort to state-of-the-art LLMs to analyze what types of privacy data are requested in natural language. Given a comprehensive list of privacy data types, our analysis model achieves 87.4% overall precision in inferring the request of a specific type of privacy data in a privacy policy. We detail our approach for this phase in Section 4.3, and evaluate our approach in Section 5.1.
Next, we explore data collection behaviors and potential data exfiltration of the collected SDK at the implementation level (shown as Phase 3). To this end, we perform an automatic static taint analysis to assess privacy data exfiltration in Android apps. Our taint analysis aims to identify the traces in the SDK implementation that constitute a complete privacy data reading and sharing process. An SDK attempts to read and share privacy data will be assessed to be prone to privacy leakage (issue type I defined in Section 2.2). We detail the design and configuration of our taint analysis in Section 4.4 and share our taint analysis results in Section 5.2. We discover the system setting injection issues from the collected SDKs and provide more details in Section 6. We also perform an additional round of evaluation and discuss it in Section 7.
In the end, we assess the status quo of privacy compliance by cross-checking the data requested in the privacy policy and the data collection behaviors of each collected SDK (Phase 4). Through this assessment, we aim to identify the excessive personal data access and permission over-claiming issues (type II and type III in Section 2.2) from the collected SDKs. We share our findings in Section 5.3. By jointly analyzing the potential data exfiltration found in SDKs’ implementation, we unveil the landscape of user privacy protection and then provide our insights and recommendations, which are detailed in Section 8.
4.2. Collecting Android SDKs
We aim to analyze data collection from developers’ disclosure and explore the data exfiltration through the lens of static taint analysis over the collected SDKs. Furthermore, we also assess privacy compliance by cross-checking the personal data requested by the developers against the actual collection behaviors found in the SDKs at the implementation level. To this end, in addition to downloading the SDK binaries (usually in AAR and JAR formats), we also collect the privacy policies and relevant metadata released by the developers, if available.

### v2

# Assessing Privacy Compliance of Android Third-Party SDKs

@@ Abstract
Third-party Software Development Kits (SDKs) are widely adopted in Android app development, to accelerate development pipelines and enhance app functionality effortlessly. However, this convenience raises substantial concerns about unauthorized access to users’ privacy-sensitive information, which could be further abused for illegitimate purposes like user tracking or monetization. Our study offers a targeted analysis of user privacy protection among Android third-party SDKs, filling a critical gap in the Android software supply chain. It focuses on two aspects of their privacy practices, including data exfiltration and behavior-policy compliance (or privacy compliance), utilizing techniques of taint analysis and large language models. It covers 158 widely-used SDKs from two key SDK release platforms, the official one and a large alternative one. From them, we identified 338 instances of privacy data exfiltration. On privacy compliance, our study reveals that more than 30% of the examined SDKs fail to provide a privacy policy to disclose their data handling practices. Among those that provide privacy policies, 37% of them over-collect user data, and 88% falsely claim access to sensitive data. We revisit the latest versions of the SDKs after 12 months. Our analysis demonstrates a persistent lack of improvement in these concerning trends. Based on our findings, we propose three actionable recommendations to mitigate the privacy leakage risks and enhance privacy protection for Android users. Our research not only serves as an urgent call for industry attention but also provides crucial insights for future regulatory interventions.
Index Terms:
mobile, privacy assessment, taint analysis.
I Introduction
@@
The advent and ubiquity of third-party Software Development Kits (SDKs) in Android app development herald a profound shift in the mobile software development paradigm, transforming efficiency and accelerating innovation. These SDKs provide pre-built reusable code that offers a broad spectrum of functionalities, from fundamental features like user authentication [1] and data encryption [2] to specialized components like analytics [3], advertisement delivery [4], and user interface enhancements [5]. The open-source nature and global reach of the Android ecosystem further fuel the proliferation of these SDKs.
@@
Integrating an SDK into an app is often as simple as declaring dependencies in the build configuration file, thus drastically reducing development time, and allowing developers to focus on app-specific features. Nowadays, the majority of Android apps incorporate multiple third-party SDKs, with an average app integrating approximately 8.6 SDKs [6, 7]. However, this widespread dependency also introduces significant security and privacy concerns. The complexity of the software supply chain and the opaque data handling practices of some SDKs can lead to privacy risks and security vulnerabilities, either due to malicious intent or unintentional oversights. These can be exploited by malevolent actors or lead to inadvertent privacy violations, as evidenced by recent high-profile incidents related to third-party SDKs [8, 9].
@@
The privacy issues caused by SDKs could be further magnified in the pre-installed apps on Android devices. These apps are bundled by device manufacturers or network carriers, and therefore, often have an extremely great user population. They come with high privileges and are typically non-removable, such that an SDK embedded by them can stealthily perform higher-privileged operations that normal apps are unable to. An illustrative example is an app named “Mobile Services Manager” pre-installed on certain Android devices, which was found to download apps without user consent due to a malicious third-party SDK embedded [10, 11].
@@
Assessing privacy practices of SDKs includes two key aspects, i.e., data exfiltration and behavior-policy compliance (or privacy compliance). The data exfiltration involves the behaviors of locating sensitive data in publicly accessible places, such as the publicly accessible directories in file systems and the Internet. The opacity of SDKs’ operations and their “black-box” nature hinders meaningful assessments, leaving both end-users and app developers in the dark and elevating risks. Dishonest or malicious SDKs may disguise as their embedding apps to extensively collect privacy-sensitive data, while the end-users are unable to distinguish which party requests their privacy data. Incidents like the X-Mode controversy [12], where user locations were tracked by over 100 apps through embedded SDKs, underscore this point and serve as a stark reminder of the challenges associated with preserving user’s privacy in third-party SDKs.
@@
On the regulatory front, data protection regulations have been put in place globally, such as the General Data Protection Regulation (GDPR) [13] and the California Consumer Privacy Act (CCPA) [14], which impose a stringent requirement on data processors. In response to this, software developers are required to provide a privacy policy to explicitly disclose what personal information is collected and how these data are handled by their software. Dissecting whether the data handling behaviors of SDKs comply with the claims in their privacy policies, which we define as behavior-policy compliance (or privacy compliance), becomes an urgent task.
@@
In this work, we conduct a large-scale study on the privacy practices of existing third-party SDKs. Our study covers 158 prevalently-installed SDKs from two mainstream release platforms, including Google Play SDK Index [15] and CAICT SDK Index [16]. The former is the official source of Android SDKs as a part of Google Play infrastructure, which lists the most widely used commercial SDKs in the Android app ecosystem [17]. The latter is claimed to be the largest release platform of mobile SDKs in China, serving as a complementary source of SDKs due to the absence of Google’s service in the Chinese market. Based on the statistical data provided by the two platforms, over 100 collected SDKs have been embedded by at least 100 apps with over 10 million installations, demonstrating the broad impact of this study.
Our study comprises two main strategies, i.e., conducting a data exfiltration assessment through static program analysis, and assessing privacy compliance through interpreting privacy policies. Specifically, we conduct a static taint analysis for tracking the flow of privacy-sensitive data within the SDK. We explore the traces that read and share privacy-sensitive data in public spaces, and accordingly investigate whether these traces can lead to potential privacy-sensitive data exfiltration. For the privacy compliance assessment, we employ a large language model (LLM) to analyze the SDKs’ privacy policies and thereby identify what data is requested, cross-check the data claimed in privacy policies with the data access behaviors caught during the taint analysis, and determine if an excessive collection or over-claiming issue exists.
@@
  • Detecting Potential Privacy Data Exfiltration on a Large Scale. We collect 158 SDKs from two mainstream repositories and leverage taint analysis to track sensitive data flow, thereby exploring their potential data exfiltration. We find 338 data exfiltration evidenced by sharing obtained privacy data into public spaces. Our work highlights the substantial privacy risks stemming from the widespread use of third-party SDKs in the Android ecosystem, unveiling complex and often opaque data-handling practices.
@@
II Backgrounds and Problem Definitions
II-A GDPR and Privacy Compliance
The General Data Protection Regulation (GDPR) [13], a landmark legislation enacted by the European Union (EU) in 2018, has profoundly transformed the landscape of data privacy laws within the region. Under GDPR, entities processing personal data, defined broadly as any data enabling the direct or indirect identification of individuals, are subjected to stringent regulations that underscore the importance of “lawfulness, fairness, and transparency”. Furthermore, GDPR provides explicit requirements for obtaining valid consent, necessitating that it is freely given, specific, informed, and unambiguous.
Privacy compliance, in the context of GDPR, necessitates a thorough examination of privacy policies to ensure adherence to the GDPR’s comprehensive principles and stipulations. It entails a meticulous inspection of data processing activities, thereby ensuring data is acquired and utilized only for explicit, legitimate purposes. Furthermore, the principle of “data minimization” obliges organizations to limit their data processing activities to the minimum necessary, thereby reinforcing the importance of data accuracy and relevance. This process also extends to an examination of storage practices, necessitating the validation that data is retained only for as long as necessary and is safeguarded against unlawful processing, inadvertent loss, or damage. Privacy compliance, therefore, functions as a crucial mechanism to enforce accountability, reinforcing the necessity for organizations to demonstrate, unequivocally, their adherence to GDPR principles.
II-B Android SDKs and Data Collection Practices
Android third-party SDKs represent a crucial component within the broader ecosystem of Android app development. These software SDKs offer pre-written code to app developers, enabling the efficient integration of different features and functions within their apps, ranging from user interface elements to complex data analytics capabilities. Data collection practices associated with Android third-party SDKs have been a subject of scrutiny within academic and industry communities. It has been observed that some SDKs, while providing utility functions, also engage in extensive data collection activities without explicit acknowledgment or consent received from the end-users. Furthermore, there are instances where data is transmitted to remote servers, raising concerns about data security and potential misuse. In light of this, there is an increasing need of a large-scale privacy assessment of SDKs.
II-C Objectives and Problem Definitions
This work aims to (1) detect potential exfiltration of users’ privacy data existing in the third-party SDKs, (2) explore the privacy compliance of developers in the SDK release, and thereby, (3) unveil the landscape of privacy protection in the Android ecosystem at the SDK level. To this end, we propose a compliance model to systematically assess the privacy practices of Android SDKs regarding personal information collection. We let $\mathbb{D}$ represent a set of diverse data considered as user privacy, and let $\mathbb{S}$ be the set of publicly released third-party SDKs. Then we define what privacy data an SDK claims to collect and actually collects as follows.
@@ II-C Objectives and Problem Definitions
We let $\mathcal{C}$ denote the compliance disclosure of an SDK $s\in\mathbb{S}$. Specifically, it defines the set of privacy data claimed to access in the privacy policy of the SDK $s$. It shall also be assumed not to leak any privacy data to the public.
@@
The compliance practice of an SDK concerns two types of operations in its implementation, i.e., collecting privacy data, and sharing the collected privacy data. We let $\mathcal{R}_{s}$ denote the scope of privacy data that would be read/collected by an SDK ${s}$ based on its implementation, in which $\mathcal{R}_{s}\subseteq\mathbb{D}$. We also let $\mathcal{U}_{s}$ denote the collected privacy data involved in any uploading/sharing operations, given that $\mathcal{U}_{s}\subseteq\mathcal{R}_{s}$. Thus, we have the compliance practice of an SDK ${s}$ as a tuple of its collecting and sharing behaviors, written as $\mathcal{P}_{s}=(\mathcal{R}_{s},\mathcal{U}_{s})$.
@@
[…]
$\exists s\in\mathbb{S},\exists d\in\mathbb{D},d\in\mathcal{U}_{s}\Rightarrow\,\,\nvdash_{\mathcal{P}}\mathcal{C}_{s}.$ (1)
@@
$\exists s\in\mathbb{S},\exists d\in\mathbb{D},(d\in\mathcal{R}_{s})\land(d\notin\mathcal{C}_{s})\Rightarrow\mathcal{R}_{s}\nvdash_{\mathcal{P}}\mathcal{C}_{s}.$ (2)
We remark that Google has not set a formal definition of “data collection” at the SDK level as what it does for third-party apps [18]. Thus, we refer to the guideline of GDPR Article 13 [13, 19] and take a conservative strategy by considering data accessed by SDKs as collected when evaluating the excessive collection, given that the data accessed by an SDK can flow into the host apps, which can then consume or send the data out.
@@
We assert that an SDK has an over-claiming issue if it is found to claim access to more types of privacy data than it actually reads. Over-claiming issues, although they may not substantially lead to privacy data exfiltration, seriously violate the data minimization principle stipulated in Article 5(1)(c) of GDPR. It is defined as follows.
$\exists s\in\mathbb{S},\exists d\in\mathbb{D},(d\notin\mathcal{R}_{s})\land(d\in\mathcal{C}_{s})\Rightarrow\mathcal{R}_{s}\nvDash_{\mathcal{P}}\mathcal{C}_{s}.$ (3)
II-D Threat Model
We consider that the exfiltration of users’ privacy data caused by malicious SDKs should follow a unified routine. To achieve such exfiltration, the malicious SDK (the attacker) gains capabilities as follows:
@@ II-D Threat Model
III Methodology
III-A Approach Overview
Figure 1: The workflow of potential data exfiltration exploration and compliance assessment at the SDK level
We propose a four-phase approach and briefly illustrate it in Figure 1. In Phase 1, we collect 158 Android SDKs, including their binary files, privacy policies, and metadata from two release platforms. After that, Phase 2 involves analyzing the privacy policies using state-of-the-art LLMs to identify the types of privacy data requested by the SDKs. In Phase 3, we perform static taint analysis on the SDKs to examine their data collection behaviors at the implementation level, aiming to identify traces that constitute a complete privacy data reading and sharing process. SDKs caught to read and share user privacy would be prone to privacy leakage (i.e., type I issues). Finally, in Phase 4, we assess privacy compliance by cross-checking the data requested in the privacy policies with the actual data collection behaviors, aiming to detect excessive collection (Type II) and over-claiming (Type III) issues.
III-B Scope of Privacy Data
Considering most privacy data stipulates an explicit permission for apps to access, we mainly resort to the list of permissions [20] from Android developer documentation to define the list of privacy data. The list is further complemented with privacy-sensitive items recognized from the privacy or security change section of each Android OS release report, for example, the clipboard and a group of user-unresettable identifiers (UUIs) are recognized in the release document of Android 10 [21] and added to the scope of this study. In addition, we also take existing literature in personal information collection [22, 23, 24, 25, 26, 27] and app analysis [28, 29, 30, 31] into consideration to complement the list of privacy data. We thus include the usage of diverse sensors (e.g., motion and light sensors) into the list because they have been widely studied in the research community. As a result, we recognize 41 types of privacy data and categorize them into five groups. These 41 data types constitute $\mathbb{D}$ defined in Section II-C and formulate the scope of our compliance assessment. The list of recognized privacy data is shown in Table I.
TABLE I: List of 41 pre-identified privacy data types and the distribution of the 741 source methods
@@ III-B Scope of Privacy Data
  • We then refine our categorization by grouping those data that can be programmatically accessed through APIs in the same or similar classes based on the Android developer’s documentation. This is to facilitate our next step’s taint analysis as we need to collect as many involved APIs as possible to be the source methods, which we will detail in Section III-E.
@@
Based on the criterion listed above, we further categorize them into five classes, namely chip, cellular & peripheral (C1), wireless communication (C2), camera, location, and other sensors, (C3), media and software specific (C4), and miscellaneous personal data (C5).
@@
Considering the request of personal data in privacy policies may not always written in a precise manner as defined in the developer’s documentation, and sometimes certain privacy data items have multiple names or colloquial phrases, we refer to relevant research [32, 33] and propose an ontology to capture the semantic relationship among the 41 types of privacy data, and consequently to guide our assessment.
First, we manually collect alternative names or colloquial phrases for each type of identified privacy data and treat them as synonyms ($\equiv$). For example, we find that the advertising ID and GAID are interchangeably used in privacy policy documents to refer to the Google Ad ID. Thus, all the synonymous terms of a privacy data type will be considered in our assessment.
@@
[…]
III-C Phase 1: Collecting Android SDKs
Our study starts by collecting as many Android SDKs as possible from mainstream sources. Since our approach involves cross-checking SDK privacy disclosures with actual data collection behaviors at the implementation level, we gather not only SDK binaries (typically in AAR and JAR formats) but also the privacy policies and relevant metadata provided by the developers, when available.